### v2

@@ 3   DID-IV in multiple time periods
We now extend the $2\times 2$ DID-IV design to multiple periods settings. Specifically, we consider the setting with the staggered adoption of the instrument across units: the exposure to the instrument occurs across units at different times and once units become exposed to the instrument, units remain exposed to that instrument afterward. We call this a staggered DID-IV design, and establish the target parameter and identifying assumptions.
@@ 3.1   Set up
We introduce the notation we use throughout section 3 to section 5. We consider a panel data setting with $T$ periods and $N$ units. For each $i\in\{1,\dots N\}$ and $t\in\{1,\dots,T\}$, let $Y_{i,t}$ denote the outcome and $D_{i,t}\in\{0,1\}$ denote the treatment status, and $Z_{i,t}\in\{0,1\}$ denote the instrument status. Let $D_{i}=(D_{i,1},\dots,D_{i,T})$ and $Z_{i}=(Z_{i,1},\dots,Z_{i,T})$ denote the path of the treatment and the instrument for unit $i$. Throughout section 3 to section 5, we assume that $\{Y_{i,t},D_{i,t},Z_{i,t}\}_{t=1}^{T}$ are independent and identically distributed (i.i.d).
@@
Assumption 7 requires that once units start exposed to the instrument, units remain exposed to that instrument afterward. In the DID literature, this assumption is made for the adoption process of the treatment and sometimes called ”staggered treatment adoption”, see, e.g., Athey and Imbens (2022), Callaway and Sant’Anna (2021) and Sun and Abraham (2021).
Exploiting variation arising from the staggered adoption of the instrument across units is pervasive in economics. For instance, Black et al. (2005) estimate the causal relationship between parents’ and children’s education by exploiting the different timing of implementation of school reforms across municipalities as an instrument for parental education.
Given Assumption 7, we can uniquely characterize one’s instrument path by the initial exposure date of the instrument, which we denote $E_{i}=\min\{t:Z_{i,t}=1\}$. If unit $i$ is not exposed to the instrument for all time periods, we define $E_{i}=\infty$. Based on the initial exposure period $E_{i}$, we can uniquely partition units into mutually exclusive and exhaustive cohorts $e$ for $e\in\{1,2,\dots,T,\infty\}$: all the units in cohort $e$ start exposed to the instrument in period $E_{i}=e$.
Similar to the two periods setting in section 2.1, we allow the general adoption process for the treatment: treatment can potentially turn on/off repeatedly over time. de Chaisemartin and D’Haultfœuille (2020) and Imai and Kim (2021) consider the same setting in the recent DID literature.
@@
Assumption 7 allows us to rewrite $D_{i,t}(z)$ by the initial exposure date $E_{i}=e$. Let $D_{i,t}^{e}$ denote the potential treatment status in period $t$ if unit $i$ is first exposed to the instrument in period $e$. Let $D_{i,t}^{\infty}$ denote the potential treatment status in period $t$ if unit $i$ is never exposed to the instrument. Hereafter, we call $D_{i,t}^{\infty}$ the ”never exposed treatment”. Since the adoption date of the instrument uniquely pins down one’s instrument path, we can write the observed treatment status $D_{i,t}$ for unit $i$ in period $t$ as
@@
We define $D_{i,t}-D_{i,t}^{\infty}$ to be the effect of an instrument on treatment for unit $i$ in period $t$, which is the difference between the observed treatment status $D_{i,t}$ to the never exposed treatment status $D_{i,t}^{\infty}$. Hereafter, we refer to $D_{i,t}-D_{i,t}^{\infty}$ as the individual exposed effect in the first stage. In the DID literature, Callaway and Sant’Anna (2021) and Sun and Abraham (2021) define the effect of a treatment on an outcome in the same fashion.
Next, we introduce the group variable which describes the type of unit $i$ in period $t$, based on the reaction of potential treatment choices in period $t$ to the instrument path $z$. Let $G_{i,e,t}\equiv(D_{i,t}^{\infty},D_{i,t}^{e})$ ($t\geq e$) be the group variable in period $t$ for unit $i$ and the initial exposure date $e$. Specifically, the first element $D_{i,t}^{\infty}$ represents the treatment status in period $t$ if unit $i$ is never exposed to the instrument $E_{i}=\infty$ and the second element $D_{i,t}^{e}$ represents the treatment status in period $t$ if unit $i$ is first exposed to the instrument at $E_{i}=e$. Following to the terminology in section 2.1, we define $G_{i,e,t}=(0,0)\equiv NT_{e,t}$ to be the never-takers, $G_{i,e,t}=(1,1)\equiv AT_{e,t}$ to be the always-takers, $G_{i,e,t}=(0,1)\equiv CM_{e,t}$ to be the compliers and $G_{i,e,t}=(1,0)\equiv DF_{e,t}$ to be the defiers in period $t$ and the initial exposure date $e$.
@@ 3.2   Target parameter in staggered DID-IV designs
Each CLATT is a natural generalization of the LATET in section 2.2 and suitable for the setting of the staggered instrument adoption. This parameter measures the treatment effects at a given relative period $l$ from the initial exposure date $E_{i}=e$, for those who belong to cohort $e$, and are the compliers $CM_{e,e+l}$, that is, who are induced to treatment by instrument in period $e+l$. Each CLATT can potentially vary across cohorts and over time because it depends on cohort $e$, relative period $l$ and the compliers $CM_{e,e+l}$. We note that the composition of the compliers can overlap over the post-exposure periods in cohort $e$. For instance, some units in cohort $e$ may belong to $CM_{e,e+l}$ for all the relative periods ($l\geq e$): they are affected by the instrument for all post-exposure periods.
This parameter would be policy-relevant if researchers leverage the policy change of their interest as an instrument for the treatment. Suppose for instance that the treatment represents participation in a job training program, the outcome represents real earnings, and the instrument represents a policy shock that encourages people to receive the treatment. In this example, if the instrument is an important policy lever for the researchers (e.g., nudge), each CLATT is of particular interest because the compliers in each relative period after a policy change represents the dynamic effect of that policy on the take up rates of the program, and this parameter can be interpreted as estimating the effect of participation in job training on real earnings among those affected in each relative period.
@@ 3.3   Identification assumptions in staggered DID-IV designs
In this section, we establish the identification assumptions in staggered DID-IV designs. These assumptions are the natural generalization of Assumptions 1-6 in $2\times 2$ DID-IV designs.
@@
Given Assumption 8 and Assumption 9, we can write the potential outcome $Y_{i,t}(d,z)$ as $Y_{i,t}(d_{t})=D_{i,t}Y_{i,t}(1)+(1-D_{i,t})Y_{i,t}(0)$. Following to section 2.3, we introduce the potential outcomes in period $t$ if unit $i$ is assigned to the instrument path $z\in\mathcal{S}(Z)$:
@@
Since the initial exposure date $E_{i}$ completely characterizes the instrument path, we can write the potential outcomes for cohort $e$ and cohort $\infty$ as $Y_{i,t}(D_{i,t}^{e})$ and $Y_{i,t}(D_{i,t}^{\infty})$ respectively. The potential outcome $Y_{i,t}(D_{i,t}^{e})$ represents the outcome status in period $t$ if unit $i$ is first exposed to the instrument in period $e$ and the potential outcome $Y_{i,t}(D_{i,t}^{\infty})$ represents the outcome status in period $t$ if unit $i$ is never exposed to the instrument. Hereafter, we refer to $Y_{i,t}(D_{i,t}^{\infty})$ as the ”never exposed outcome”.
@@
This assumption requires that the instrument path affects the treatment adoption behavior in a monotone way for all relative periods after the initial exposure date $E_{i}=e$. We recall that we define $D_{i,t}-D_{i,t}^{\infty}$ to be the effect of an instrument on treatment for unit $i$ in period $t$. Assumption 10 requires that the individual exposed effect in the first stage should be non-negative (or non-positive) during the periods after the initial exposure to the instrument for all $i$. This assumption implies that the group variable $G_{i,e,t}\equiv(D_{i,t}^{\infty},D_{i,t}^{e})$ can take three values with non-zero probability for all $e$ and all $t\geq e$. Hereafter, we consider the type of the monotonicity assumption that rules out the existence of the defiers $DF_{e,t}$ for all $t\geq e$ in any cohort $e$.
@@
Assumption 12 is a parallel trends assumption in the treatment in multiple periods and multiple cohorts. This assumption requires that the trends of the treatment across cohorts would have followed the same path, on average, if there is no exposure to the instrument. Assumption 12 is analogous to that of Callaway and Sant’Anna (2021) and Sun and Abraham (2021) in DID designs: these studies impose the same type of parallel trends assumption on untreated outcomes with settings of multiple periods and multiple cohorts.
@@
Assumption 13 is a parallel trends assumption in the outcome with settings of multiple periods and multiple cohorts. This assumption requires that the expectation of the never exposed outcome across cohorts would have followed the same evolution if the assignment of the instrument had not occurred. From the discussion in section 2.4, we can interpret that this assumption requires the same expected time gain across cohorts and over time: the effects of time on the outcome through treatment are the same on average across cohorts and over time.
@@ 4   Extensions
This section contains extensions to non-binary, ordered treatments, triple DID-IV designs, and repeated cross sections. It also includes the case when we introduce the treatment path in potential outcomes in $2\times 2$ DID-IV designs. For more details and the proofs in this section, see Appendix B.
@@
Up to now, we have considered only the case of a binary treatment. However the same idea can be applied when treatment takes a finite number of ordered values, that is, we have $D_{i,t}\in\{0,1,\dots,J\}$. When we have only two periods and treatment is non-binary, our target parameter is the average causal response on the treated (ACRT) defined below.
@@
The average causal response on the treated (ACRT) is a weighted average of the effects of a unit increase in the treatment on the outcome, for those who belong to an exposed group and are induced to increase the treatment in period $t=1$ by instrument. The ACRT is similar to the average causal response (ACR) considered in Angrist and Imbens (1995), but the difference here is that each weight $w_{j}$ and the associated causal parameter in this parameter are conditional on $E_{i}=1$.
Theorem 3 below shows that if we have a non-binary, ordered treatment, the Wald-DID estimand is equal to the average causal response on the treated (ACRT) under the same assumptions in Theorem 1.
@@
If we have a non-binary, ordered treatment in staggered DID-IV designs, our target parameter is the cohort specific average causal response on the treated (CACRT), which is the natural generalization of the ACRT. The identifying assumptions are the same in section 3.3.
@@
In section 2.1, we impose Assumption 1 (No carryover assumption) on potential outcomes $Y_{i,t}(d,z)$ as we consider the non-staggered adoption of the treatment. We can extend our results to the case where we introduce the treatment path in potential outcomes with binary treatment. In this case, our target parameter slightly changes with the cumbersome notation, formally defined below.
@@
In Appendix B.3, we show that the similar assumptions as in Assumptions 1-6 in section 2.3 are sufficient for the Wald-DID to capture the LATET’.
@@ 5   Estimation and Inference
When researchers leverage variation in the timing of policy adoption across units as an instrument for the treatment and implicitly rely on a staggered DID-IV design, they usually implement this design via TWFEIV regressions (e.g., Johnson and Jackson (2019), Lundborg et al. (2014), Black et al. (2005), Akerman et al. (2015), and Bhuller et al. (2013)).
Unfortunately, however, recent research has shown that in more than two periods, the TWFEIV estimand potentially fails to summarize the treatment effects under staggered DID-IV designs in the presence of heterogeneous treatment effects. Specifically, Miyaji (2024) shows that under staggered DID-IV designs, the TWFEIV estimand is a weighted average of all possible CLATTs, but some weights can be negative [^6] See also the discussions in de Chaisemartin and D’Haultfœuille (2020), who decompose the numerator and denominator in the TWFEIV estimand separately, and point out the issue of interpreting this estimand causally.. Miyaji (2024) further shows that the negative weight problem arises if the effect of the instrument on the treatment or the outcome is not stable over time.
In this section, we propose an alternative estimation method in staggered DID-IV designs that is robust to treatment effect heterogeneity. Specifically, we first propose a simple regression-based method for estimating each CLATT. Following to Callaway and Sant’Anna (2021), we then propose a weighting scheme to construct the summary causal parameters from each CLATT. At the end of this section, we also discuss the pretrends tests for checking the plausibility of parallel trends assumption in the treatment and the outcome.
@@ 5.1   Stacked two stage least squares regression
In this section, we propose a regression-based method to consistently estimate our target parameter in staggered DID-IV designs. Recall that if we have a binary treatment, our target parameter is the $CLATT_{e,e+l}$ for each cohort $e$ and a relative period $l\geq 0$:
@@
Researchers should carefully specify the choice set $U$ in practice. If there exists a never exposed cohort $E_{i}=\infty$, we can set $U=\{\infty\}$ and estimate $CLATT_{e,e+l}$ for all cohorts ($e\in\{2,\dots,T\}$) and all relative time periods ($l\geq 0$). If there exists no never exposed cohort, we can set $U=\max\{E_{i}\}$, i.e., the last exposed cohort. In this case, we can estimate $CLATT_{e,e+l}$ for cohort $e\in\{2,\dots,\max\{E_{i}\}-1\}$ and relative time period $l\in\{0,\dots,\max\{E_{i}\}-1-e\}$ because every unit will be exposed to the instrument after time $t=\max\{E_{i}\}$. We note that the units in an already exposed cohort $e=1$ are not included in any data sets because we should include the pre-exposed period $e-1$ for estimating each $CLATT_{e,e+l}$.
@@
Here, $E_{N}[\cdot]$ is the sample analog of the conditional expectation. We note that the IV estimator $\widehat{CLATT}_{e,e+l}$ is the Wald-DID estimator for each $CLATT_{e,e+l}$, where the choice set $U$ is an unexposed group and time $t=e-1$ is a pre-exposed period. From the estimation procedure above, we call this a stacked two-stage least squares (STS) estimator.
Our STS estimator is related to the DID estimators in staggered DID designs recently proposed by Callaway and Sant’Anna (2021) and Sun and Abraham (2021). Specifically, the DID estimator of the treatment and the outcome in our STS estimator corresponds to that of Sun and Abraham (2021), and coincides with that of Callaway and Sant’Anna (2021) for the case when we have no covariates and use never treated units as a control group. Their proposed methods avoid the issue of two-way fixed effects estimators in staggered DID designs, and our STS estimator avoids the issue of two-way fixed effects instrumental variable estimators in staggered DID-IV designs.
@@
Suppose Assumptions 7-13 hold. Then, the STS estimator $\widehat{CLATT}_{e,e+l}$ is consistent and asymptotically normal:
@@
Our estimation procedure is identical if we have a non-binary, ordered treatment or use repeated cross section data. In Appendix C, we present the influence function of our SLS estimator in repeated cross section settings.
@@
where $\mathbf{1}_{A}$ is indicator function and takes one if $A$ is true. In Appendix C, we also present the influence function of our triple DID-IV estimator $\beta_{IV}^{e,l}$.

### v4

@@ 3   DID-IV in multiple time periods
We now extend the $2\times 2$ DID-IV design to multiple period settings with the staggered adoption of the instrument across units. In reality, many studies exploit variation in the timing of policy adoption across units as an instrument for treatment in more than two periods (e.g. Black et al. (2005), Bhuller et al. (2013), Lundborg et al. (2014), and Meghir et al. (2018)). We refer to the underlying identification strategy as a staggered DID-IV design, and establish the target parameter and identifying assumptions.
@@ 3.1   Set up
We introduce the notation we use throughout Section 3 to Section 5. We consider a panel data setting with $T$ periods and $N$ units. For each $i\in\{1,\dots N\}$ and $t\in\{1,\dots,T\}$, let $Y_{i,t}$ denote the outcome, $D_{i,t}\in\{0,1\}$ denote the treatment status, and $Z_{i,t}\in\{0,1\}$ denote the instrument status. Let $D_{i}=(D_{i,1},\dots,D_{i,T})$ and $Z_{i}=(Z_{i,1},\dots,Z_{i,T})$ denote the path of the treatment and the instrument for unit $i$, respectively. Throughout Section 3 to Section 5, we assume that $\{Y_{i,t},D_{i,t},Z_{i,t}\}_{t=1}^{T}$ are independent and identically distributed (i.i.d).
@@
Assumption 7 requires that once units start getting exposed to the instrument, units remain exposed to that instrument. In the DID literature, several studies make a similar assumption for the treatment and call it the ”staggered treatment adoption” (e.g. Athey and Imbens (2022); Callaway and Sant’Anna (2021); and Sun and Abraham (2021)).
Given Assumption 7, one can uniquely characterize one’s instrument path by the initial exposure date of the instrument, denoted as $E_{i}=\min\{t:Z_{i,t}=1\}$. If unit $i$ is not exposed to the instrument for all time periods, we define $E_{i}=\infty$. Based on the initial exposure period $E_{i}$, one can uniquely partition units into mutually exclusive and exhaustive cohorts $e$ for $e\in\{1,2,\dots,T,\infty\}$: all the units in cohort $e$ start being exposed to the instrument in period $E_{i}=e$.
Similar to the two periods setting in Subsection 2.1, we allow the general adoption process for the treatment: the treatment can potentially turn on/off repeatedly over time. de Chaisemartin and D’Haultfœuille (2020) and Imai and Kim (2021) also consider the same setting in the recent DID literature.
@@
Assumption 7 allows us to rewrite $D_{i,t}(z)$ by the initial exposure date $E_{i}=e$. Let $D_{i,t}^{e}$ denote the potential treatment status in period $t$ if unit $i$ is first exposed to the instrument in period $e$. Let $D_{i,t}^{\infty}$ denote the potential treatment status in period $t$ if unit $i$ is never exposed to the instrument. We call $D_{i,t}^{\infty}$ the ”never exposed treatment”. Since the adoption date of the instrument uniquely pins down one’s instrument path, we can write the observed treatment status $D_{i,t}$ for unit $i$ in period $t$ as
@@
We define $D_{i,t}-D_{i,t}^{\infty}$ to be the effect of an instrument on treatment for unit $i$ in period $t$, which is the difference between the observed treatment status $D_{i,t}$ to the never exposed treatment status $D_{i,t}^{\infty}$. We refer to $D_{i,t}-D_{i,t}^{\infty}$ as the individual exposed effect in the first stage. In the DID literature, Callaway and Sant’Anna (2021) and Sun and Abraham (2021) define the effect of a treatment on an outcome in the same fashion.
Next, we introduce the group variable that describes the type of unit $i$ in period $t$, based on the reaction of potential treatment choices in period $t$ to the instrument path $z$. Let $G_{i,e,t}\equiv(D_{i,t}^{\infty},D_{i,t}^{e})$ ($t\geq e$) be the group variable in period $t$ for unit $i$ and the initial exposure date $e$. The first element $D_{i,t}^{\infty}$ represents the treatment status in period $t$ if unit $i$ is never exposed to the instrument $E_{i}=\infty$ and the second element $D_{i,t}^{e}$ represents the treatment status in period $t$ if unit $i$ is first exposed to the instrument at $E_{i}=e$. Following the terminology in section 2.1, we define $G_{i,e,t}=(0,0)\equiv NT_{e,t}$ as the never-takers, $G_{i,e,t}=(1,1)\equiv AT_{e,t}$ as the always-takers, $G_{i,e,t}=(0,1)\equiv CM_{e,t}$ as the compliers and $G_{i,e,t}=(1,0)\equiv DF_{e,t}$ as the defiers in period $t$ and the initial exposure date $e$.
@@ 3.2   Target parameter in staggered DID-IV designs
Each CLATT is a natural generalization of the LATET in Subsection 2.2 and suitable for the setting of the staggered instrument adoption. This parameter measures the treatment effects at a given relative period $l$ from the initial exposure date $E_{i}=e$, for those who belong to cohort $e$, and are the compliers $CM_{e,e+l}$ who are induced to treatment by instrument in period $e+l$. Each CLATT can potentially vary across cohorts and over time because it depends on cohort $e$, relative period $l$, and the compliers $CM_{e,e+l}$. Note that the composition of the compliers can overlap over the post-exposure periods in cohort $e$. For instance, some units in cohort $e$ may belong to $CM_{e,e+l}$ for all the relative periods ($l\geq e$): they are affected by the instrument for all post-exposure periods.
This parameter would be policy-relevant if the instrument is based on the policy change of interest to the researcher. For example, suppose that the treatment represents participation in a job training program, the outcome represents real earnings, and the instrument represents a policy shock that encourages people to receive the treatment. In this setting, if the instrument is an important policy lever for the researchers (e.g., nudge), each CLATT is of particular interest; the share of compliers in each relative period after the policy shock represents the dynamic effect of that policy change on the take-up rate of the program, and the corresponding CLATT can be interpreted as estimating the effect of participation in job training on real earnings among those affected.
@@ 3.3   Identification assumptions in staggered DID-IV designs
In this subsection, we establish the identification assumptions in staggered DID-IV designs. These assumptions are the natural generalization of Assumptions 1-6 in $2\times 2$ DID-IV designs.
@@
Given Assumption 8 and Assumption 9, we can write the potential outcome $Y_{i,t}(d,z)$ as $Y_{i,t}(d_{t})=D_{i,t}Y_{i,t}(1)+(1-D_{i,t})Y_{i,t}(0)$. Following Subsection 2.3, we introduce the potential outcomes in period $t$ if unit $i$ is assigned to the instrument path $z\in\mathcal{S}(Z)$.
@@
Since the initial exposure date $E_{i}$ completely characterizes the instrument path, we can write the potential outcomes for cohort $e$ and cohort $\infty$ as $Y_{i,t}(D_{i,t}^{e})$ and $Y_{i,t}(D_{i,t}^{\infty})$, respectively. The potential outcome $Y_{i,t}(D_{i,t}^{e})$ represents the outcome status in period $t$ if unit $i$ is first exposed to the instrument in period $e$ and the potential outcome $Y_{i,t}(D_{i,t}^{\infty})$ represents the outcome status in period $t$ if unit $i$ is never exposed. We refer to $Y_{i,t}(D_{i,t}^{\infty})$ as the ”never exposed outcome”.
@@
This assumption requires that the instrument path affects the treatment adoption behavior in a monotone way for all relative periods after the initial exposure date $E_{i}=e$. Recall that we define $D_{i,t}-D_{i,t}^{\infty}$ to be the effect of an instrument on treatment for unit $i$ in period $t$. Assumption 10 requires that the individual exposed effect in the first stage should be non-negative (or non-positive) during the periods after the initial exposure to the instrument for all $i$. This assumption implies that the group variable $G_{i,e,t}\equiv(D_{i,t}^{\infty},D_{i,t}^{e})$ can take three values with non-zero probability for all $e$ and all $t\geq e$. Hereafter, we consider the type of the monotonicity assumption that rules out the existence of the defiers $DF_{e,t}$ for all $t\geq e$ in any cohort $e$.
@@
Assumption 12 is a parallel trends assumption in the treatment in multiple periods and cohorts. It requires that the trends of the treatment across cohorts would have followed the same path, on average, if there was no exposure to the instrument. Assumption 12 is analogous to that of Callaway and Sant’Anna (2021) and Sun and Abraham (2021) in DID designs; these studies impose the same type of parallel trends assumption on untreated outcomes in settings of multiple periods and cohorts.
@@
Assumption 13 is a parallel trends assumption in the outcome with settings of multiple periods and cohorts. It requires that the expectation of the never exposed outcome across cohorts would have followed the same evolution if the instrument had not been assigned. From the discussion in Subsection 2.4, this assumption requires the same expected time gain across cohorts and over time: the effects of time on the outcome through treatment are the same on average across cohorts and over time.
@@ 4   Extensions
This section contains extensions to non-binary, ordered treatments, triple DID-IV designs, and repeated cross sections. It includes the case in which the treatment path was introduced in potential outcomes in $2\times 2$ DID-IV designs. For more details and proofs, see Appendix B.
@@
Up to now, we have considered only the case of a binary treatment. However the same idea can be applied when treatment takes a finite number of ordered values; that is, we have $D_{i,t}\in\{0,1,\dots,J\}$. When we have only two periods and treatment is non-binary, our target parameter is the average causal response on the treated (ACRT) defined below.
@@
The ACRT is a weighted average of the effects of an increase of one unit in the treatment on the outcome, for those who belong to an exposed group and are induced to increase the treatment in period $t=1$ by instrument. The ACRT is similar to the average causal response (ACR) considered in Angrist and Imbens (1995), with the difference that each weight $w_{j}$ and the associated causal parameter in this parameter are conditional on $E_{i}=1$.
Theorem 3 below shows that if we have a non-binary, ordered treatment, the Wald-DID estimand is equal to the ACRT under the same assumptions in Theorem 1.
@@
With a non-binary, ordered treatment in staggered DID-IV designs, our target parameter is the cohort specific average causal response on the treated (CACRT), a natural generalization of the ACRT. The identifying assumptions are the same as in Subsection 3.3.
@@
In Subsection 2.1, we impose Assumption 1 (no carryover assumption) on potential outcomes $Y_{i,t}(d,z)$ as we consider the non-staggered adoption of the treatment. We can extend our results to the case where we introduce the treatment path in potential outcomes with binary treatment. In this case, our target parameter changes slightly with some complex notation.
@@
Appendix B.3 shows that the similar assumptions to Assumptions 1-6 in Subsection 2.3 are sufficient for the Wald-DID to capture the LATET’.
@@ 5   Estimation and inference
When researchers leverage variation in policy adoption timing across units as an instrument for treatment and implicitly rely on a staggered DID-IV design, they usually implement this design via TWFEIV regressions (e.g., Johnson and Jackson (2019), Lundborg et al. (2014), Black et al. (2005), Akerman et al. (2015), and Bhuller et al. (2013)).
The companion paper (Miyaji (2024)), however, shows that in more than two periods, the TWFEIV estimand potentially fails to summarize the treatment effects under staggered DID-IV designs in the presence of heterogeneous treatment effects. Specifically, it shows that under staggered DID-IV designs, the TWFEIV estimand is a weighted average of all possible CLATTs, but some weights can be negative if the effect of the instrument on the treatment or the outcome is not stable over time. [^7] See also the discussions in de Chaisemartin and D’Haultfœuille (2020), who decompose the numerator and denominator in the TWFEIV estimand separately, and point out the issue of interpreting this estimand causally.
In this section, we propose an alternative estimation method in staggered DID-IV designs that is robust to treatment effect heterogeneity. First, we propose a simple regression-based method for estimating each CLATT. Following Callaway and Sant’Anna (2021), we then propose a weighting scheme to construct the summary causal parameters from each CLATT. At the end of this section, we also discuss the pre-trend tests for checking the plausibility of parallel trends assumption in the treatment and the outcome in DID-IV designs.
@@ 5.1   Stacked two stage least squares regression
We propose a regression-based method to consistently estimate our target parameter in staggered DID-IV designs. If we have a binary treatment, the target parameter is the $CLATT_{e,e+l}$ for each cohort $e$ and a relative period $l\geq 0$:
@@
Researchers should carefully specify the choice set $U$ in practice. If a never exposed cohort $E_{i}=\infty$ exists, we can set $U=\{\infty\}$ and estimate $CLATT_{e,e+l}$ for all cohorts ($e\in\{2,\dots,T\}$) and all relative time periods ($l\geq 0$). If no never exposed cohorts exist, we can set $U=\max\{E_{i}\}$, the last exposed cohort, and can estimate $CLATT_{e,e+l}$ for cohort $e\in\{2,\dots,\max\{E_{i}\}-1\}$ and relative time period $l\in\{0,\dots,\max\{E_{i}\}-1-e\}$ because every unit will be exposed to the instrument after time $t=\max\{E_{i}\}$. Note that the units in an already exposed cohort $e=1$ are not included in any data sets because the pre-exposed period $e-1$ should be included for estimating each $CLATT_{e,e+l}$.
@@
Here, $E_{N}[\cdot]$ is the sample analog of the conditional expectation. Note that the IV estimator $\widehat{CLATT}_{e,e+l}$ is the Wald-DID estimator for each $CLATT_{e,e+l}$, where the choice set $U$ is an unexposed group and time $t=e-1$ is a pre-exposed period. From the estimation procedure, we call this a stacked two-stage least squares (STS) estimator.
Our STS estimator is related to the DID estimators in staggered DID designs proposed by Callaway and Sant’Anna (2021) and Sun and Abraham (2021). The DID estimator of the treatment and the outcome in our STS estimator corresponds to that of Sun and Abraham (2021), and coincides with that of Callaway and Sant’Anna (2021) for the case when no covariates exist and never treated units are used as a control group. Their proposed methods avoid the issue of TWFE estimators in staggered DID designs, and our STS estimator avoids the issue of TWFEIV estimators in staggered DID-IV designs.
@@
Suppose Assumptions 7-13 hold. Then, the STS estimator $\widehat{CLATT}_{e,e+l}$ is consistent and asymptotically normal.
@@
Our estimation procedure is identical if we have a non-binary, ordered treatment or use repeated cross section data. Appendix C presents the influence function of the SLS estimator in repeated cross section settings.
@@
where $\mathbf{1}_{A}$ is the indicator function and takes one if $A$ is true. Appendix C presents the influence function of the triple DID-IV estimator $\beta_{IV}^{e,l}$.